the vertex of this parabola is at (2,-1). when the y-value is 0, the x-value is 5. What is the coefficient of the squared term in the parabola's equation?
Answer by dfrazzetto(283)

Vertex form:
y = a(x-2)^2 -1
solve for a:
0 = a(5-2)^2 -1
1 = a(3)^2
1 = 9a
a = 1/9